# A Survey on Autonomous Driving Datasets: Data Statistic, Annotation, and Outlook

### v1

@@ V-B2 Multi-Task Datasets
INTERACTION. The INTERACTION [20] dataset covers diverse, complex, and critical driving scenes, coupled with its comprehensive semantic map, allowing it to be a versatile platform for a multitude of tasks, such as motion prediction, imitation learning, and validation of decision and planning. Its inclusion of different countries further improves the robustness of analyzing the driving behavior across different cultures, which is critical for the global development of autonomous driving.
BLVD. The BLVD [106] benchmark facilitates tasks such as dynamic 4D (3D+temporal) tracking, 5D (4D+interactive) interactive event recognition, and intention prediction, which are essential for a deeper understanding of traffic scenes. BLVD offers around 120K frames from different traffic scenes, comprising object density (low and high) and illumination conditions (daytime and nighttime). These frames are fully annotated with a large number of 3D labels incorporating vehicles, pedestrians, and riders.
rounD. The rounD dataset presented by [107] is pivotal for scenario classification, road user behavior prediction, and driver modeling because of the large number of collections of road user trajectories at roundabouts. The dataset utilizes a drone equipped with a 4K resolution camera to collect over six hours of video, recording more than 13K road users. The broad recorded traffic situations and the high-quality recordings make rounD an essential dataset in autonomous driving, facilitating the study of naturalistic driving behaviors in public traffic.
Lyft Level 5. Lyft Level 5 [108] represents one of the most extensive autonomous driving datasets for motion prediction to date, with over 1,000 hours of data. It encompasses 17,000 25-second long scenes, a high-definition semantic map with over 15,000 human annotations, 8,500 lane segments, and a high-resolution aerial image of the area. It supports multiple tasks like motion forecasting, motion planning, and simulation. The numerous multimodal data with detailed annotations make the Lyft Level 5 dataset a vital benchmark for prediction and planning.
@@
SceNDD. SceNDD [110] introduces real driving scenarios that feature diverse trajectories and driving behaviors for developing efficient motion planning and path-following algorithms. It is also adaptable to different configurations of the ego-vehicle and contains a time horizon of prediction that can be broken down into timestamps for detailed analysis. In conclusion, the SceNDD dataset is a significant addition to autonomous driving prediction and planning research.
DeepAccident. The synthetic dataset, DeepAccident [12], is the first work that provides direct and explainable safety evaluation metrics for autonomous vehicles. The extensive dataset with 57K annotated frames and 285K annotated samples supports end-to-end motion and accident prediction, which is vital for improving the predictive capabilities of autonomous driving systems in avoiding collisions and ensuring safety. Moreover, this multimodal dataset is versatile for various V2X-based perception tasks, such as 3D object detection, tracking, and bird’s-eye-view (BEV) semantic segmentation.
Talk2Bev. The innovative dataset, Talk2BEV [111], promotes the shift from traditional autonomous driving tasks to incorporate large vision-language models with BEV maps in the context of autonomous driving. Talk2BEV utilizes recent advances in vision-language models, allowing for a more flexible and comprehensive understanding of road scenarios. This dataset comprises over 20,000 diverse question categories, all human-annotated and derived from [22]. The proposed Talk2BEV-Bench benchmark can be leveraged across multiple tasks, covering decision-making, visual and spatial reasoning, and intent prediction.
V2X-Seq (Forecasting). The trajectory forecasting dataset is a substantial part of the real-world dataset V2X-Seq [50], containing about 80K infrastructure-view and 80K vehicle-view scenarios, and a further 50K cooperative-view scenarios. This diversity of sensing domains creates a more holistic view of the traffic environment, exhibiting huge potentiality for research and analysis on vehicle-infrastructure-cooperative (VIC) trajectory prediction.
@@ V-C End-to-End Datasets
DDD17. The DDD17 [113] dataset is notable for its use of event-based cameras, which provide a concurrent stream of standard active pixel sensor (APS) images and dynamic vision sensors (DVS) temporal contrast events, offering a unique blend of visual data. Additionally, DDD17 captures diverse driving scenarios, including highway and city driving and various weather conditions, thus providing exhaustive and realistic data for training and testing end-to-end autonomous driving algorithms.
Other datasets summarized in this survey are shown in Tab. IV Tab. V, Tab. VI.
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
nighttime drive [116] 2018 35K $\times$ onboard SS real
[…]
A*3D [119] 2020 39,179 point cloud frames $\times$ onboard (3D) OD real
PreSIL [8] 2019 50K frames $\checkmark$ onboard OD, 3D SS synthetic
H3D [120] 2019 27,721 frames $\checkmark$ onboard (3D) OD, MOT real
ROAD [121] 2021 122K frames $\checkmark$ onboard OD, SS real
CADC [123] 2020 56K $\times$ onboard (3D) OD, OT real
PandaSet [128] 2021 8,240 frames $\checkmark$ onboard (3D) OD, SS, OT real
TUD-Brussels [132] 2009 1,600 frames $\times$ onboard (3D) OD real
VEIS [133] 2018 61,305 frames $\times$ onboard OD, SS synthetic
IDDA [136] 2020 1M frames $\times$ onboard segmentation synthetic
TJ4RadSet [137] 2022 7,757 frames $\checkmark$ onboard OD, OT real
TME Motorway [138] 2012 30K frames $\checkmark$ onboard OD, OT real
CARRADA [140] 2020 7,193 radar frames $\checkmark$ onboard SS real
SODA10M [141] 2021 20K labeled images $\checkmark$ onboard OD real
Robo3D [142] 2023 476K frames $\checkmark$ onboard (3D) OD, 3D SS real
CODA [145] 2022 1,500 frames $\times$ onboard corner case detection real
PixSet [146] 2021 29K point cloud frames $\times$ onboard (3D) OD real
K-Radar [147] 2023 35K radar frames $\times$ onboard 3D OD, OT real
VIL-100 [149] 2021 10K frames $\times$ onboard lane detection real
EuroCity Persons [150] 2018 47,300 frames $\times$ onboard OD real
WADS [153] 2021 1K point cloud frames $\checkmark$ onboard SS real
NightCity [154] 2020 4,297 frames $\times$ onboard nighttime SS real
LLMAS [156] 2019 100K frames $\checkmark$ onboard lane detection real
Toronto-3D [158] 2020 4 scenarios $\times$ onboard 3D SS real
RANUS [163] 2018 4K frames $\times$ onboard SS, scene understanding real
AUTOCASTSIM [165] 2022 52K frames $\checkmark$ V2X (3D) OD, OT, SS real
Mcity [166] 2019 1,7500 frames $\checkmark$ onboard SS real
TRoM [169] 2017 712 frames $\times$ onboard road marking detection real
Cirrus [170] 2020 6,285 frames $\checkmark$ onboard (3D) OD real
Multifog KITTI [173] 2021 15K frames $\times$ onboard 3D OD synthetic
UNDD [174] 2019 7.2K frames $\checkmark$ onboard SS real
RaidaR [176] 2021 58,542 rainy street scenes $\times$ onboard SS real
K-Lane [177] 2022 15,382 frames $\times$ onboard lane detection real
aiMotive [178] 2023 26,583 frames $\checkmark$ onboard 3D OD, MOT real
SAP [179] 2016 19K frames $\times$ drone OD, OT real
Astyx [180] 2019 500 radar frames $\times$ onboard 3D OD real
Comap [181] 2021 4,391 frames $\checkmark$ V2X 3D OD synthetic
WIBAM [185] 2021 33,092 frames $\checkmark$ V2X 3D OD real
MIT DriveSeg [186] 2019 5K frames $\times$ onboard SS real
TABLE IV: Autonomous driving dataset 1
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
DRIV100 [189] 2021 100K frames $\times$ onboard domain adaptation SS real
NEOLIX [190] 2021 30K frames $\checkmark$ onboard 3D OD, OT real
IPS3000+ [192] 2021 14,198 frames $\checkmark$ V2X 3D OD real
TUM-Traffic [193] 2023 4.8K frames $\checkmark$ V2X 3D OD, OT real
Amodal Cityscapes [194] 2022 5K frames $\times$ onboard amodal SS real
CrashD [204] 2022 15,340 scenes $\times$ onboard 3D OD synthetic
MIT DSS [186] 2020 10K video frames $\checkmark$ onboard SS real
CPIS [83] 2020 10K frames $\times$ V2X cooperative 3D OD synthetic
DDAD [209] 2020 21,200 frames $\times$ onboard depth estimation real
Ua-detrac [210] 2015 140K frames $\checkmark$ V2X OD, MOT real
NCLT [211] 2015 34.9 hours $\checkmark$ onboard odometry real
NVSEC [212] 2018 around 28 km distances $\checkmark$ others SLAM real
MSLU [213] 2013 36.8 km distances $\checkmark$ onboard SLAM real
A2D2 [25] 2020 433,833 frames $\checkmark$ onboard 3D OD, SS, SLAM real
Complex Urban [217] 2017 around 190km paths onboard SLAM real
Syncapes [218] 2018 25K images $\times$ onboard OD, SS synthetic
Paris-rue-Madame [222] 2014 643 objects $\times$ onboard OD, SS real
CADP [226] 2018 1,416 scenes $\times$ V2X traffic accident analysis real
NYC3DCars [229] 2013 2K images $\times$ onboard OD real
RUGD [230] 2019 37K frames $\checkmark$ others SS real
EU LTD [231] 2020 around 37 hours $\checkmark$ onboard odometry real
LLamas [156] 2019 100,042 images $\times$ onboard lane marker detection, lane segmentation real
CCD [234] 2020 4.5K videos $\checkmark$ onboard accident prediction real
 4Seasons [238] 2020 350km recordings $\checkmark$ onboard SLAM real
Boxy [241] 2019 200K frames $\checkmark$ onboard OD real
AMUSE [242] 2013 117,440 frames $\checkmark$ onboard SLAM real
Brno Urban [243] 2019 375.7 km $\checkmark$ onboard recognition real
DurLAR [245] 2021 100K frames $\checkmark$ onboard depth estimation real
PVDN [248] 2021 59,746 frames $\checkmark$ onboard nighttime OD, OT real
RoadSaW [250] 2022 720K scenes $\checkmark$ onboard Road surface and wetness estimation real
ASAP [252] 2022 1.2M images $\checkmark$ onboard online 3D OD real
Brain4Cars [255] 2015 2M frames $\checkmark$ onboard Maneuver Anticipation real
NEXET [256] 2017 91,190 frames $\times$ onboard OD real
DIML [257] 2017 470 videos $\checkmark$ onboard lane detection real
TABLE V: Autonomous driving dataset 2
Dataset Year Size Temporal Sensing domain Tasks Real/Synthetic
VLMV [261] 2020 900 frames $\checkmark$ V2X lane merge real
PREVENTION [263] 2019 356 mins $\checkmark$ onboard onboard trajectory and intention prediction real
TABLE VI: Autonomous driving dataset 3
@@ VI Annotations process
VI-A How annotations are created
Different AD tasks require specific types of annotation. For example, object detection requires bounding box labels of instance, segmentation is based on pixel- or point-level annotations, and continually labeled trajectory is critical for trajectory prediction. On the other hand, as shown in Fig. 11, the annotation pipeline can be categorized into three types: manual annotation, semi-automatic annotation, and fully automatic annotation. We detail the labeling approaches for different types of annotation in this section.
@@ VI-A How annotations are created
Many studies have proposed fully or semi-automatic annotation methods to improve annotation efficiency. [264] presented a fully automatic annotation approach based on weakly supervised learning to segment proposed drivable paths in images. [265] is a semi-automatic annotation method utilizing the objectness priors to generate segmentation masks. After that, [266] offered a semi-automatic method considering 20 classes. Polygon-RNN++[267] presented an interactive segmentation annotation tool following the idea of [268]. Instead of using image information to generate pixel-level labels, [269] explored transferring 3D information into 2D image domains to generate semantic segmentation annotations. For labeling 3D data, [270] proposed an image-assisted annotation pipeline. [271] leveraged active learning to select a few points and to form a minimal training set to avoid labeling the whole point cloud scenes. [272] introduced an efficient labeling framework with semi/weakly supervised learning to label outdoor point clouds.
Labelme [273] is a prior tool focusing on labeling images for object detection. However, generating bounding boxes by professional annotators faces the same issue as manual segmentation annotation. Wang et al. [274] presented a semi-automatic video labeling tool based on the open-source video annotation system VATIC [^7] https://www.cs.columbia.edu/~vondrick/vatic/. [275] is another automatic video annotation tool for AD scenes. Compared to daytime annotation, process bounding box annotations in the nighttime is more challenging. [276] introduced a semi-automatic approach leveraging the trajectory to solve this problem.
@@
As one of the prior works in trajectory annotation, [280] online generated actions for maneuvers and were annotated into the trajectory. [281] consists of a crowd-sourcing step followed by a precise process of expert aggregation. [282] developed an active learning framework to annotate driving trajectory. Precisely anticipating movement patterns of pedestrians is critical for driving safety. Styles et al. [283] introduced a scalable machine annotation scheme for pedestrian trajectory annotations without human effort.
@@
The generated synthetic scenarios are designed to mimic real-world conditions, including multiple objects, various landscapes, weather conditions, and lighting variations. To achieve this goal, some researchers utilized the Grand Theft Auto 5 (GTA5) game engine to build datasets [4, 5]. [284] presented a real-time system based on multiple games to generate annotations for various AD tasks. Instead of applying game videos, SHIFT [10], CAOS [7], and V2XSet [11] were created based on the CARLA [112] simulator. Compared to [11], V2X-Sim [81] studied employing multiple simulators [112, 285] to generate dataset for V2X perception tasks. CODD [205] further exploited using [112] to generate 3D LiDAR point clouds for cooperative driving. Other works [3, 9, 2, 133] leveraged the Unity development platform [286] to generate synthetic datasets.
VI-B The quality of annotations
Existing supervised learning-based AD algorithms build upon numerous labeled data. However, training on low-quality annotations can negatively affect the safety and reliability of autonomous vehicles. Therefore, ensuring the quality of annotations is fundamental for improving accuracy while driving in complex real-world environments.
@@ VI-B The quality of annotations
According to the study [287], the annotation quality is affected by several factors, such as consistency, correctness, precision, and validation. Consistency is the foremost criterion in evaluating annotation quality. It involves maintaining uniformity across the entire dataset and is crucial for avoiding confusion in models trained on this data. For example, if a particular type of vehicle is labeled as a car, it should be consistently annotated the same in all other instances. Annotation precision is another vital indicator, which refers to whether the labels match the actual state of the objects or scenarios. In contrast, correctness highlights that annotated data are appropriate and relevant for the dataset purpose and annotation guidelines. After annotation, it is essential to validate the annotated data to ensure its accuracy and completion. The process can be done through manual review by experts or algorithms. Validation helps effectively prevent issues in datasets before they infect the performance of autonomous vehicles, decreasing potential safety risks. [288] presented a data-agnostic validation method for expert annotated datasets.
A failure case of annotation from KITTI [13] is shown in Fig. 12. We illustrate the ground truth bounding boxes (blue) in the corresponding image and LiDAR point cloud. On the left side of the image, the annotation of a car (circled in red) is inaccurate because it does not contain the whole object car. Additionally, two cars (highlighted by the green cuboid) are not annotated, even though the camera and LiDAR capture them clearly.
@@ VII-A Worldwide Distribution
We demonstrate an overview of the global distribution of 191 autonomous driving datasets in Fig. 13. The chart indicates that the USA is at the forefront with 40 datasets (21% share), underscoring its leadership in the autonomous driving domain. Germany accounts for 24 datasets, reflecting its robust automotive industry and influence on the advancement of autonomous technology. China is closely followed with 16 datasets, suggesting the interest and investment in this field of China. Another notable point is that there are 11 datasets collected worldwide and 24 from the European region (except Germany). This diverse regional distribution enhances the robustness of the collected data and highlights the international efforts and collaborations in the research community and industry.
On the other hand, although smaller segments represent a range of other countries, including Canada, Korea, the UK, Japan, and Singapore, these countries are developed countries with solid technological backgrounds and accumulation—an extreme regional bias reflected by this statistic. The dominance of the USA, western Europe, and East Asia leads to the bias where autonomous driving systems are overfitted to environmental conditions typical of these regions. This bias could result in autonomous vehicles’ failure to perform under varied or unseen regions and cases. Hence, introducing data from a more expansive array of countries and regions, such as Africa, can assist in the comprehensive development of autonomous vehicles.
Furthermore, 35 synthetic datasets generated by simulators like CARLA [112] take up 18.32% percent. Due to the limitation of recording from real-world driving environments, these synthetic datasets overcome such drawbacks and are critical for exploiting more robust and reliable driving systems.

### v2

@@ V-B2 Multi-Task Datasets
INTERACTION. The INTERACTION [20] dataset is a versatile platform that offers diverse, complex, and critical driving scenes, along with a comprehensive semantic map, making it suitable for a multitude of tasks, such as motion prediction, imitation learning, and validation of decision and planning. Its inclusion of different countries and continents further improves the robustness of analyzing the driving behavior across different cultures. A potential shortcoming of the dataset is that the impact of environmental conditions has not been explicitly addressed.
rounD. The rounD dataset presented by [122] is pivotal for scenario classification, road user behavior prediction, and driver modeling. It provides a large number of road user trajectories at roundabouts. The dataset utilizes a drone equipped with a 4K resolution camera to collect over six hours of video, recording more than 13K road users. The broad recorded traffic situations and the high-quality recordings make rounD an essential dataset in autonomous driving, facilitating the study of naturalistic driving behaviors in public traffic. Yet, similar to all datasets only collected under clear weather conditions, developing models capable of performing under challenging situations could be restricted.
Lyft Level 5. Lyft Level 5 [125] represents one of the most extensive autonomous driving datasets for motion prediction to date, with over 1,000 hours of data. It encompasses 17,000 25-second long scenes, a high-definition semantic map with over 15,000 human annotations, 8,500 lane segments, and a high-resolution aerial image of the area. It supports multiple tasks like motion forecasting, motion planning, and simulation. The numerous multimodal data with detailed annotations make it a vital benchmark for prediction and planning. The dataset shows limitations in accurately representing scenes with uncommon traffic conditions or rare pedestrian behaviors.
@@
DeepAccident. The synthetic dataset, DeepAccident [12] is the first work that provides direct and explainable safety evaluation metrics for autonomous vehicles. The extensive dataset with 57K annotated frames and 285K annotated samples supports end-to-end motion and accident prediction, which is vital in avoiding collisions and ensuring safety. Moreover, this multimodal dataset is versatile for various V2X-based perception tasks, such as 3D object detection, tracking, and BEV semantic segmentation. The various environmental conditions also enhance the generalizability of this dataset. Due to the domain adaptation issue, the performance of models trained on DeepAccident in real driving situations needs further study.
@@ V-C End-to-end Datasets
DDD17. The DDD17 [129] dataset is notable for using event-based cameras. The dataset provides a concurrent stream of standard active pixel sensor (APS) images and dynamic vision sensors (DVS) temporal contrast events, offering a unique blend of visual data. Additionally, DDD17 captures diverse driving scenarios, including highway and city driving and various weather conditions, thus providing exhaustive and realistic data for training and testing end-to-end autonomous driving algorithms.
@@ VI Annotations process
VI-A Annotation Generation
Different AD tasks require specific types of annotation. For example, object detection requires bounding box labels of instance, segmentation is based on pixel- or point-level annotations, and continually labeled trajectory is critical for trajectory prediction. On the other hand, as shown in Fig. 11, the annotation pipeline can be categorized into three types: manual annotation, semi-automatic annotation, and fully automatic annotation. In this section, we detail the labeling approaches for different annotation types.
Labelme [130] is a prior tool focusing on labeling images for object detection. However, generating bounding boxes by professional annotators faces the same issue as manual segmentation annotation. Wang et al. [131] presented a semi-automatic video labeling tool based on the open-source video annotation system VATIC[132]. Manikandan et al. [133] propose another automatic video annotation tool for AD scenes. Process bounding box annotations in the nighttime is more challenging than daytime annotation. Schorkhuber et al. [134] introduced a semi-automatic approach leveraging the trajectory to solve this problem.
@@ VI-A Annotation Generation
Many studies have proposed fully or semi-automatic annotation methods to improve annotation efficiency. Barnes et al. [139] presented a fully automatic annotation approach based on weakly supervised learning to segment proposed drivable paths in images. The semi-automatic annotation method [140] utilizes the objectness priors to generate segmentation masks. After that, [141] offered a semi-automatic method considering 20 classes. Polygon-RNN++[142] presents an interactive segmentation annotation tool following the idea of [143]. Instead of using image information to generate pixel-level labels, [137] explores transferring 3D information into 2D image domains to generate semantic segmentation annotations. For labeling 3D data, [144] proposes an image-assisted annotation pipeline. Luo et al. [145] leverage active learning to select a few points and to form a minimal training set to avoid labeling the whole point cloud scenes. Liu et al. [146] introduce an efficient labeling framework with semi/weakly supervised learning to label outdoor point clouds.
@@
As one of the prior works in trajectory annotation, [147] online generates actions for maneuvers and is annotated into the trajectory. The annotation approach [148] consists of a crowd-sourcing step followed by a precise process of expert aggregation. Jarl et al. [149] develop an active learning framework to annotate driving trajectory. Precisely anticipating movement patterns of pedestrians is critical for driving safety. Styles et al. [150] introduce a scalable machine annotation scheme for pedestrian trajectory annotations without human effort.
@@
The generated synthetic scenarios are designed to mimic real-world conditions, including multiple objects, various landscapes, weather conditions, and lighting variations. Selecting a suitable simulation tool is crucial to achieve this goal. Torcs [151] and DeepDriving [152] are two prior works for simulating autonomous vehicles while lacking multi-modality information and other types of objects like pedestrians. Recently, the open source simulators, such as CARLA [128], SUMO [153], and AirSim [154], have been widely used in data generation. They provide transparent platforms where researchers and developers can freely access and modify the source code, tailoring the simulator to specific requirements. In contrast, because of non-open sources, commercial simulation tools like NVIDIA’s Drive Constellation [155] can make it difficult for users to create special driving environments. Rather than professional simulators, the game engines, including Grand Theft Auto 5 (GTA5) and Unity [156], are also popular for creating synthetic autonomous driving data.
Specifically, some researchers utilize the GTA5 game engine to build datasets [4] and [5]. Krahenbuhl et al. [157] present a real-time system based on multiple games to generate annotations for various AD tasks. Instead of applying game videos, SHIFT [10], CAOS [7], FedBEVT [158], and V2XSet [11] are created based on the CARLA simulator. Compared to [11], V2X-Sim [89] studies employing multiple simulators [128, 153] to generate dataset for V2X perception tasks. CODD [159] further exploits using [128] to generate 3D LiDAR point clouds for cooperative driving. Other works [3, 9, 2, 160] leverage the Unity development platform to generate synthetic datasets.
VI-B Quality of Annotations
Supervised learning-based autonomous driving (AD) algorithms rely heavily on extensive, well-labeled datasets. High-quality datasets ensure that these systems can accurately perceive and interpret complex driving environments, enhancing safety and reliability on the road. This in turn fosters user trust, a crucial factor for the widespread adoption of autonomous vehicles. Conversely, poor dataset quality can result in system errors and safety risks, undermining user confidence, hindering acceptance, and failing to meet the criteria for trustworthy AI as discussed by [161]. Therefore, ensuring the quality of annotations is fundamental for improving accuracy while driving in complex real-world environments. It is even more important to fine-tune the data quality and use active learning methods for dataset curation to get robust performance results on a test set than fine-tuning the model architecture [162].
@@ VI-B Quality of Annotations
According to the study [163], the annotation quality is affected by several factors, such as consistency, correctness, precision, and validation. Consistency is the foremost criterion in evaluating annotation quality. It involves maintaining uniformity across the entire dataset and is crucial for avoiding confusion in models trained on this data. For example, if a particular type of vehicle is labeled as a car, it should be consistently annotated the same in all other instances. Annotation precision is another vital indicator, which refers to whether the labels match the actual state of the objects or scenarios. In contrast, correctness highlights that annotated data are appropriate and relevant for the purpose of the dataset and annotation guidelines. After annotation, it is essential to validate the annotated data to ensure its accuracy and completion. The process can be done through manual review by experts or algorithms. Validation helps effectively prevent issues in datasets before they infect the performance of autonomous vehicles, decreasing potential safety risks. Inel et al. [164] presented a data-agnostic validation method for expert annotated datasets.
A failure case of annotation from KITTI [13] is shown in Fig. 12. We illustrate the ground truth bounding boxes (blue) in the corresponding image and LiDAR point cloud. On the left side of the image, the annotation of a car (circled in red) is inaccurate because it does not contain the whole object car. Additionally, two cars (highlighted by the green rectangle) are not annotated, even though the camera and LiDAR capture them clearly. Furthermore, datasets like the IPS300+ [165] contain many labeled objects within a frame (319.84 labels per frame), but on the other hand, the annotations are of bad quality. Many large datasets like the Pandaset [166], Oxford [167], CADC [168], nuScenes [22], and Lyft Level 5 [22] were labeled by specialized labeling companies like Scale AI and provide high-quality annotations. Labeling the nuScenes dataset took about 7,937 hours and cost 100k USD. Another way to label datasets is to use custom labeling tools like 3D BAT [138, 60], which was used to create the TUMTraf dataset. The Waymo and KITTI datasets were both labeled with custom labeling tools. V2V4Real [87] has used the SUSTechPoints [169] labeling tool to generate the dataset.
@@ VII-A Worldwide Distribution
We demonstrate an overview of the global distribution of 191 autonomous driving datasets in Fig. 13. The chart indicates that the USA is at the forefront with 40 datasets (21% share), underscoring its leadership in the autonomous driving domain. Germany accounts for 24 datasets, and China closely follows with 16 datasets. On the other hand, developed countries, including Canada, Korea, the UK, Japan, and Singapore, cover the rest smaller segments. Although 11 datasets are collected worldwide and 24 are from the European region (except Germany), all these countries or regions are considered high-income areas. The dominance of the USA, western Europe, and East Asia reflects an extremely unbalanced development of autonomous driving around the world.
Specifically, one of the most classic datasets, KITTI [13], is collected in the metropolitan area of Karlsruhe, Germany. In comparison, both Waymo [23] and Argoverse 2 [28] datasets are collected from a wide variety of places, including six different cities in the USA individually. Apolloscapes [17] and DAIR-V2X [27] are recorded in China. Instead of collecting data solely within a country, nuScenes [22] is established based on the data from America (Boston) and Singapore, both known for complex and highly challenging traffic situations. Other well-known autonomous driving datasets [24, 88, 95, 18, 21, 118] are collected in those previously mentioned countries. It is noted that thanks to the geographic diversity, [22, 23] have been widely used in transfer learning to verify the generalizability of autonomous driving algorithms.
Moreover, autonomous driving faces unique challenges across different geographical regions. However, relying solely on data from a single source can introduce bias that could result in autonomous vehicles’ failure to perform under varied or unseen regions and cases. For instance, the diversity and quantity of electric scooters in China far exceed those in Germany, meaning an algorithm trained exclusively on German data may struggle to accurately recognize objects in China. Hence, Recording data from different continents and countries can assist in addressing unique challenges caused by geographical locations. This diverse regional distribution enhances the robustness of the collected data and highlights the international efforts and collaborations in the research community and industry.
Furthermore, 35 synthetic datasets generated by simulators like CARLA [128] take up 18.32% percent. Due to the limitation of recording from real-world driving environments, these synthetic datasets overcome such drawbacks and are critical for exploiting more robust and reliable driving systems. However, the domain adaptation from synthetic data to real-world data is still a challenging research topic, which limits a broader application of synthetic data and relevant simulators.